OpenStudy (anonymous):

write each precent as a fraction im simplest form and a decimal 0.01 and 0.02 write each precent as a fraction im simplest form and a decimal 0.01 and 0.02 @Mathematics

OpenStudy (anonymous):

in*

OpenStudy (anonymous):

0.01% = 1/1000 0.0001 .02% = 1/5000 .0002

OpenStudy (karatechopper):

1% 1/100 and 2% 1/50